How many gram formula weights of H2SO4?


1
Expert's answer
2021-02-26T06:00:33-0500

Solution:

H2SO4 = sulfuric acid

The formula weight of H2SO4 = 2 × Ar(H) + Ar(S) + 4 × Ar(O)

The formula weight of H2SO4 = 2×1.00784 + 32.065 + 4×15.999 = 98.077 (g/mol)

Hence, 1 mol of sulfuric acid (H2SO4) has a weight of 98.077 g.


Answer: The formula weight of H2SO4 is 98.077 grams.